OpenStudy (anonymous):

evaluate the logarithmic expression log5 1/125? show work

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

\[\log_{a}a \]=1

OpenStudy (gorv):

when base and number are same log =1

OpenStudy (gorv):

\[\log_{5}1-\log_{5}125 \]

OpenStudy (gorv):

log1=0 0-log5 (5*5*5)=-log5*5^3=-3*log5 5=-3*1=-3

OpenStudy (anonymous):

why is it -3 instead of 3? @gorv

OpenStudy (gorv):

u knw??? log(a/b)=loga -logb

OpenStudy (gorv):

@ralph123

OpenStudy (anonymous):

okay thanks @gorv